# Re-anchoring Quantum Monte Carlo with Tensor-Train Sketching

### v4

@@ 3. Proposed Algorithm
[…]
@@
A Hamiltonian $H$, trial wavefunction $\Psi_{\mathrm{tr},0}$, step size $\Delta\tau$, number of iterations $M_{\text{out}}$ and $M_{\text{in}}$, sketching functions $S$ and $T$.
An ensemble of walkers $\widehat{\Psi}^{(M_{\text{in}}M_{\text{out}})}=\sum_{k}w^{(M_{\text{in}}M_{\text{out}})}_{k}\frac{\phi^{(M_{\text{in}}M_{\text{out}})}_{k}}{O_{\mathrm{tr}}(\phi^{(M_{\text{in}}M_{\text{out}})}_{k})}$, a wavefunction $\Psi_{\mathrm{tr},M_{\text{out}}}$.
Let $\widehat{\Psi}^{(0)}=\sum_{k}\Psi_{\mathrm{tr,0}}$ (i.e. $w_{0,k}=1,\phi_{0,k}=\Psi_{\mathrm{tr}}$ which implies $O_{\mathrm{tr}}(\phi_{0,k})=1$).
@@
  • $\widehat{\Psi}^{(M_{\text{in}}n_{\text{out}}+M_{\text{in}})}$ = cp-AFQMC($H,\widehat{\Psi}^{(M_{\text{in}}n_{\text{out}})},\Psi_{\mathrm{tr},n_{\text{out}}},\Delta\tau,M_{\text{in}}$).
  • $\Psi_{\mathrm{tr},n_{\text{out}}+1}$ = TT-sketching($\widehat{\Psi}^{(M_{\text{in}}n_{\text{out}}+M_{\text{in}})}$,$S$,$T$).
@@ 4.1. Variance and bias per step of cp-AFQMC
In this section, we provide an explanation on how importance sampling controls the variance of the overlaps in one step of cp-AFQMC, and discuss the systematic error caused by constrained-path approximation. We first consider a different representation of the walkers for the convenience of the analysis, where we define the following to be the walker that we keep track of:
(17) $\Phi_{k}^{(n)}=w_{k}^{(n)}\frac{\phi_{k}^{(n)}}{O_{\mathrm{tr}}(\phi^{(n)}_{k})}.$
Under this notation, based on (8) (line 2), the update rule of the walker in cp-AFQMC algorithm becomes
$\Phi_{k}^{(n+1)}=\widetilde{B}_{k}^{(n)}(\vec{x})\Phi_{k}^{(n)},$
where
(18) $\widetilde{B}_{k}^{(n)}(\vec{x}):=\frac{P(\vec{x})}{\widetilde{P}_{k}^{(n)}(\vec{x})}B(\vec{x}),\quad\vec{x}\sim\widetilde{P}_{k}^{(n)}.$
The probability distribution of importance sampling is defined in (6). It is assumed that $\widetilde{B}_{k}^{(n)}(\vec{x})$ is only well-defined when $\widetilde{P}_{k}^{(n)}(\vec{x})>0$. Again based on (8), we hope
@@
in a certain sense. We first state a simple observation regarding the update $\widetilde{B}_{k}^{(n)}(\vec{x})\Phi_{k}^{(n)}$:
@@
Suppose we have a current walker $\Phi_{k}^{(n)}$ in the form of (17), the overlap of the walker at the $(n+1)$-th step is then
$\braket{\Psi_{\mathrm{tr}},\widetilde{B}_{k}^{(n)}(\vec{x})\Phi_{k}^{(n)}}=\left\langle\Psi_{\mathrm{tr}},\frac{P(\vec{x})}{\widetilde{P}_{k}^{(n)}(\vec{x})}B(\vec{x})\Phi_{k}^{(n)}\right\rangle=\braket{\Psi_{\mathrm{tr}},\Phi_{k}^{(n)}}\mathcal{N}_{k}^{(n)}.$
@@
Lemma 1 tells us that the variance of $\braket{\Psi_{\mathrm{tr}},\widetilde{B}_{k}^{(n)}(\vec{x})\Phi_{k}^{(n)}}$ is reduced by importance sampling, in fact, to 00. We then study the value of $\braket{\Psi_{\mathrm{tr}},\widetilde{B}_{k}^{(n)}(\vec{x})\Phi_{k}^{(n)}}$, as a consequence of Lemma 1, in the following two lemmas.
@@
When $\braket{\Psi_{\mathrm{tr}},\widetilde{B}_{k}^{(n)}(\vec{x})\Phi_{k}^{(n)}}>0$, then $\widetilde{P}_{k}^{(n)}(\vec{x})>0$ for all $\vec{x}$ (see (6)). In this case, $\widetilde{B}_{k}^{(n)}(\vec{x})$ is well-defined for every $\vec{x}$ as the definition of $\widetilde{B}_{k}^{(n)}(\vec{x})$ involves division by $\widetilde{P}_{k}^{(n)}(\vec{x})$. Then,
$\braket{\Psi_{\mathrm{tr}},\widetilde{B}_{k}^{(n)}(\vec{x})\Phi_{k}^{(n)}}=\braket{\Psi_{\mathrm{tr}},\sum_{\vec{x}}\widetilde{P}_{k}^{(n)}(\vec{x})\widetilde{B}_{k}^{(n)}(\vec{x})\Phi_{k}^{(n)}}=\braket{\Psi_{\mathrm{tr}},\sum_{\vec{x}}P(\vec{x})B(\vec{x})\Phi_{k}^{(n)}}=\braket{\Psi_{\mathrm{tr}},e^{-\Delta\tau H}\Phi_{k}^{(n)}},$
where the first equality is due to Lemma 1 and the fact that $\widetilde{B}_{k}^{(n)}(\vec{x})$ can be defined for every $\vec{x}$, and the second and third equality follows from (6) and (7) respectively. ∎
It might be the case where $\widetilde{P}_{k}^{(n)}(\vec{x})=0$ for some auxiliary fields $\vec{x}$. In this situation, some of the $\vec{x}$ values cannot be reached. Hence, a systematic error on $\braket{\Psi_{\mathrm{tr}},\widetilde{B}_{k}^{(n)}(\vec{x})\Phi_{k}^{(n)}}$ is introduced, as explained in the following.
@@
$\braket{\Psi_{\mathrm{tr}},\widetilde{B}_{k}^{(n)}(\vec{x})\Phi_{k}^{(n)}}-\braket{\Psi_{\mathrm{tr}},e^{-\Delta\tau H}\Phi_{k}^{(n)}}=-\sum_{\vec{x}:\widetilde{P}_{k}^{(n)}(\vec{x})=0}P(\vec{x})\braket{\Psi_{\mathrm{tr}},B(\vec{x})\Phi_{k}^{(n)}}\geq 0.$
@@
When $\braket{\Psi_{\mathrm{tr}},B(\vec{x})\Phi_{k}^{(n)}}\leq 0$ for some $\vec{x}$, there are values of $\vec{x}$ where $\widetilde{P}_{k}^{(n)}(\vec{x})=0$. In this case, (7) can be expressed more explicitly as
$\begin{split}e^{-\Delta\tau H}&=\sum_{\vec{x}:\widetilde{P}_{k}^{(n)}(\vec{x})=0}P(\vec{x})B(\vec{x})+\sum_{\vec{x}:\widetilde{P}_{k}^{(n)}(\vec{x})>0}P(\vec{x})B(\vec{x})\\ &=\sum_{\vec{x}:\widetilde{P}_{k}^{(n)}(\vec{x})=0}P(\vec{x})B(\vec{x})+\sum_{\vec{x}:\widetilde{P}_{k}^{(n)}(\vec{x})>0}\widetilde{P}_{k}^{(n)}(\vec{x})\widetilde{B}_{k}^{(n)}(\vec{x}).\end{split}$
@@
$\begin{split}&\quad\braket{\Psi_{\mathrm{tr}},e^{-\Delta\tau H}\Phi_{k}^{(n)}}\\ &=\braket{\Psi_{\mathrm{tr}},\sum_{\vec{x}:\widetilde{P}_{k}^{(n)}(\vec{x})=0}P(\vec{x})B(\vec{x})\Phi_{k}^{(n)}+\sum_{\vec{x}:\widetilde{P}_{k}^{(n)}(x)>0}\widetilde{P}_{k}^{(n)}(\vec{x})\widetilde{B}_{k}^{(n)}(\vec{x})\Phi_{k}^{(n)}}\\ &=\sum_{\vec{x}:\widetilde{P}_{k}^{(n)}(\vec{x})=0}P(\vec{x})\braket{\Psi_{\mathrm{tr}},B(\vec{x})\Phi_{k}^{(n)}}+\sum_{\vec{x}:\widetilde{P}_{k}^{(n)}(\vec{x})>0}\widetilde{P}_{k}^{(n)}(\vec{x})\braket{\Psi_{\mathrm{tr}},\widetilde{B}_{k}^{(n)}(\vec{x})\Phi_{k}^{(n)}}\\ &=\sum_{\vec{x}:\widetilde{P}_{k}^{(n)}(\vec{x})=0}P(\vec{x})\braket{\Psi_{\mathrm{tr}},B(\vec{x})\Phi_{k}^{(n)}}+\braket{\Psi_{\mathrm{tr}},\widetilde{B}_{k}^{(n)}(\vec{x})\Phi_{k}^{(n)}}\end{split}$
@@
In this case, the actual inner product $\braket{\Psi_{\mathrm{tr}},\widetilde{B}_{k}^{(n)}(\vec{x})\Phi_{k}^{(n)}}$ is larger than $\braket{\Psi_{\mathrm{tr}},e^{-\Delta\tau H}\Phi_{k}^{(n)}}$, thus a bias towards the trial function $\Psi_{\mathrm{tr}}$ is introduced, which causes the sign problem.
@@ 4.2. Convergence analysis
In this section, we present how importance sampling with a trial wavefunction facilitates the convergence of energy. We focus on the case where there is no sign problem as in Lemma 2. We want to show that in such a situation, having a trial wavefunction that is closer to the ground-state wavefunction gives a smaller variance on the energy estimator (Section 4.2.1). We further show in Section 4.2.2 that unlike typical power method, we cannot expect that the walkers converge to the ground-state wavefunction as the variance of the individual walkers, controlled primarily by the chosen form of $B(\vec{x})$, is too large.
@@ 4.2.1. Convergence analysis of the energy
In this section, we analyze how the mixed energy estimator converges to the ground-state energy in one step of cp-AFQMC. We want to show that the variance of cp-AFQMC is controlled by $\left|c_{\perp}/c_{0}\right|$. When $\left|c_{\perp}/c_{0}\right|$ gets smaller, the variance of the energy gets smaller, i.e. the convergence of cp-AFQMC behaves like a vanilla power iteration. In an ideal case where $\left|c_{\perp}/c_{0}\right|=0$, i.e. the trial function is exactly the ground-state wavefunction $\Psi_{\mathrm{tr}}=\Psi_{0}$, there is a zero-variance principle. Thus, we are motivated to improve the trial wavefunction in Algorithm 3 to reduce the variance in ground-state energy estimation.
@@
Proof of both Lemma 4 and Theorem 5 can be found in Appendix B. We notice that the energy convergence consists of a deterministic power iteration part and a fluctuating part that is roughly $a_{0}\|H-E_{0}I\|_{2}\tan\theta_{k}^{(n)}=\left|c_{\perp}/c_{0}\right|\|H-E_{0}I\|_{2}\tan\theta_{k}^{(n)}$ (when $a_{0}$ is small). Here $\|H-E_{0}I\|_{2}$ resembles the energy scale, and $\tan\theta_{k}^{(n)}$ is a random variable that captures the deviation of a random walker from the ground state. While currently, we do not have a characterization of the variance of $\tan\theta_{k}^{(n)}$, if we assume that it has a bounded variance of $O(1)$, when $a_{0}$ gets small, i.e. the trial function gets closer to the ground-state wavefunction, the fluctuation term can be made small. In an ideal case where $a_{0}=0$ so that the trial function is exactly the ground-state wavefunction $\Psi_{\mathrm{tr}}=\Psi_{0}$, there is no fluctuation in the energy estimation. As a matter of fact, we have
@@ 4.2.2. Convergence analysis of the walkers
where $\mathcal{P}_{\perp}$ is the orthogonal projector that projects to the space spanned by $\{\Psi_{1},\ldots\Psi_{2^{d}-1}\}$. We want to show in the following that $\tan\theta_{k}^{(n+1)}(\vec{x})<\tan\theta_{k}^{(n)}$ up to some noise terms. However, the noise terms will not vanish even for very good trial wavefunction.
@@
Proof of Theorem 6 can be found in Appendix B. As we can see, even when $a_{0}=0$, i.e. the trial wavefunction is exactly the ground-state wavefunction, there is still an error term $O\left(\frac{\|\widetilde{B}_{k}^{(n)}(\vec{x})-e^{-\Delta\tau H}\|}{|\braket{\Psi_{0},\Phi_{k}^{(n)}}|}\right)$. Thus, the individual walker is not guaranteed to get closer to the ground state. In fact, in general, one needs an exponential number of walkers to get $\frac{\|\widetilde{B}_{k}^{(n)}(\vec{x})-e^{-\Delta\tau H}\|}{|\braket{\Psi_{0},\Phi_{k}^{(n)}}|}<1$. This is due to the fact that, generally, one needs to sample an exponential number of $\widetilde{B}_{k}^{(n)}(\vec{x})$ to approximate $\exp(-\Delta\tau H)$. In other words, when sampling a multi-dimensional function, we can evaluate various properties with a polynomial cost, but to map out the function to pre-specified details by relying on the sample positions can require exponential cost.

### v5

@@ 3. Proposed Algorithm
$\widehat{\Psi}^{(0)}$$\Psi_{\mathrm{tr,0}}$$\widehat{\Psi}^{(M_{\text{in}})}$$\Psi_{\mathrm{tr,1}}$$\widehat{\Psi}^{(2M_{\text{in}})}$$\Psi_{\mathrm{tr,2}}\cdotsrandom$ walkcp approx.sketchingrandom walkcp approx.sketching
@@
A Hamiltonian $H$ and a corresponding decomposition $e^{-\Delta\tau H}=\sum_{\vec{x}}P(\vec{x})B(\vec{x})$, initial trial wavefunction $\Psi_{\mathrm{tr},0}$, step size $\Delta\tau$, number of iterations $M_{\text{out}}$ and $M_{\text{in}}$.
An ensemble of walkers $\widehat{\Psi}^{(M_{\text{in}}M_{\text{out}})}=\sum_{k}\Phi^{(M_{\text{in}}M_{\text{out}})}_{k}$, a wavefunction $\Psi_{\mathrm{tr},M_{\text{out}}}$.
Let $\widehat{\Psi}^{(0)}=\sum_{k}\Psi_{\mathrm{tr,0}}$.
@@
[…]
  • Apply TT-sketching on $\widehat{\Psi}^{(M_{\text{in}}n_{\text{out}}+M_{\text{in}})}$ and get a new trial wavefunction $\Psi_{\mathrm{tr},n_{\text{out}}+1}$.
@@ 4.1. Variance and bias per step of cp-AFQMC
In this section, we provide an explanation on how importance sampling controls the variance of the overlaps in one step of cp-AFQMC, and discuss the systematic error caused by constrained-path approximation. For convenience, we rewrite the walker propagation (9) as
$\Phi_{k}^{(n+1)}:=\widetilde{B}_{k}^{(n)}(\vec{x}_{k}^{(n)})\Phi_{k}^{(n)},\vec{x}_{k}^{(n)}\sim Q_{k}^{(n)},$
where we denote the operator
(11) $\widetilde{B}_{k}^{(n)}(\vec{x})=\frac{P(\vec{x})}{Q_{k}^{(n)}(\vec{x})}B(\vec{x}),$
which is only well-defined when $Q_{k}^{(n)}(\vec{x})>0$. Based on (8), we hope
@@
in a certain sense. We firstly state an observation regarding the update $\widetilde{B}_{k}^{(n)}(\vec{x})\Phi_{k}^{(n)}$.
@@
Suppose we have a current walker $\Phi_{k}^{(n)}$, the overlap of the walker at the next step is then
$\braket{\Psi_{\mathrm{tr}},\widetilde{B}_{k}^{(n)}(\vec{x})\Phi_{k}^{(n)}}=\left\langle\Psi_{\mathrm{tr}},\frac{P(\vec{x})}{Q_{k}^{(n)}(\vec{x})}B(\vec{x})\Phi_{k}^{(n)}\right\rangle=\braket{\Psi_{\mathrm{tr}},\Phi_{k}^{(n)}}\mathcal{N}_{k}^{(n)}.$
@@
Lemma 1 tells us that the variance of $\braket{\Psi_{\mathrm{tr}},\widetilde{B}_{k}^{(n)}(\vec{x})\Phi_{k}^{(n)}}$ is reduced by importance sampling, in fact, to 0. We then study the value of $\braket{\Psi_{\mathrm{tr}},\widetilde{B}_{k}^{(n)}(\vec{x})\Phi_{k}^{(n)}}$, as a consequence of Lemma 1, in the following two lemmas.
@@
When $\braket{\Psi_{\mathrm{tr}},\widetilde{B}_{k}^{(n)}(\vec{x})\Phi_{k}^{(n)}}>0$, then $Q_{k}^{(n)}(\vec{x})>0$ for all $\vec{x}$ (see (6)). In this case, $\widetilde{B}_{k}^{(n)}(\vec{x})$ is well-defined for every $\vec{x}$. Then
$\braket{\Psi_{\mathrm{tr}},\widetilde{B}_{k}^{(n)}(\vec{x})\Phi_{k}^{(n)}}=\braket{\Psi_{\mathrm{tr}},\sum_{\vec{x}}Q_{k}^{(n)}(\vec{x})\widetilde{B}_{k}^{(n)}(\vec{x})\Phi_{k}^{(n)}}=\braket{\Psi_{\mathrm{tr}},\sum_{\vec{x}}P(\vec{x})B(\vec{x})\Phi_{k}^{(n)}}=\braket{\Psi_{\mathrm{tr}},e^{-\Delta\tau H}\Phi_{k}^{(n)}},$
where the first equality is due to Lemma 1 and the fact that $\widetilde{B}_{k}^{(n)}(\vec{x})$ can be defined for every $\vec{x}$, and the second and third equality follow from (11) and (4), respectively. ∎
It might be the case where $Q_{k}^{(n)}(\vec{x})=0$ for some auxiliary fields $\vec{x}$. In this situation, some of the $\vec{x}$ values cannot be reached. Hence, a systematic error on $\braket{\Psi_{\mathrm{tr}},\widetilde{B}_{k}^{(n)}(\vec{x})\Phi_{k}^{(n)}}$ is introduced, as explained in the following.
@@
$\braket{\Psi_{\mathrm{tr}},\widetilde{B}_{k}^{(n)}(\vec{x})\Phi_{k}^{(n)}}-\braket{\Psi_{\mathrm{tr}},e^{-\Delta\tau H}\Phi_{k}^{(n)}}=-\sum_{\vec{x}:Q_{k}^{(n)}(\vec{x})=0}P(\vec{x})\braket{\Psi_{\mathrm{tr}},B(\vec{x})\Phi_{k}^{(n)}}\geq 0.$
@@
When $\braket{\Psi_{\mathrm{tr}},B(\vec{x})\Phi_{k}^{(n)}}\leq 0$ for some $\vec{x}$, there are values of $\vec{x}$ such that $Q_{k}^{(n)}(\vec{x})=0$. In this case, (7) can be expressed more explicitly as
$\begin{split}e^{-\Delta\tau H}&=\sum_{\vec{x}:Q_{k}^{(n)}(\vec{x})=0}P(\vec{x})B(\vec{x})+\sum_{\vec{x}:Q_{k}^{(n)}(\vec{x})>0}P(\vec{x})B(\vec{x})\\ &=\sum_{\vec{x}:Q_{k}^{(n)}(\vec{x})=0}P(\vec{x})B(\vec{x})+\sum_{\vec{x}:Q_{k}^{(n)}(\vec{x})>0}Q_{k}^{(n)}(\vec{x})\widetilde{B}_{k}^{(n)}(\vec{x}).\end{split}$
@@
$\begin{split}&\quad\braket{\Psi_{\mathrm{tr}},e^{-\Delta\tau H}\Phi_{k}^{(n)}}\\ &=\braket{\Psi_{\mathrm{tr}},\sum_{\vec{x}:Q_{k}^{(n)}(\vec{x})=0}P(\vec{x})B(\vec{x})\Phi_{k}^{(n)}+\sum_{\vec{x}:Q_{k}^{(n)}(x)>0}Q_{k}^{(n)}(\vec{x})\widetilde{B}_{k}^{(n)}(\vec{x})\Phi_{k}^{(n)}}\\ &=\sum_{\vec{x}:Q_{k}^{(n)}(\vec{x})=0}P(\vec{x})\braket{\Psi_{\mathrm{tr}},B(\vec{x})\Phi_{k}^{(n)}}+\sum_{\vec{x}:Q_{k}^{(n)}(\vec{x})>0}Q_{k}^{(n)}(\vec{x})\braket{\Psi_{\mathrm{tr}},\widetilde{B}_{k}^{(n)}(\vec{x})\Phi_{k}^{(n)}}\\ &=\sum_{\vec{x}:Q_{k}^{(n)}(\vec{x})=0}P(\vec{x})\braket{\Psi_{\mathrm{tr}},B(\vec{x})\Phi_{k}^{(n)}}+\braket{\Psi_{\mathrm{tr}},\widetilde{B}_{k}^{(n)}(\vec{x})\Phi_{k}^{(n)}}\end{split}$
@@
In this case, the actual inner product $\braket{\Psi_{\mathrm{tr}},\widetilde{B}_{k}^{(n)}(\vec{x})\Phi_{k}^{(n)}}$ is larger than the expected value $\braket{\Psi_{\mathrm{tr}},e^{-\Delta\tau H}\Phi_{k}^{(n)}}$, thus a bias towards the trial function $\Psi_{\mathrm{tr}}$ is introduced, which causes the sign problem.
@@ 4.2. Convergence analysis
In this section, we present how importance sampling with a trial wavefunction facilitates the convergence of energy. We focus on the case where there is no sign problem as in Lemma 2. We want to show that in this, having a trial wavefunction that is closer to the ground-state wavefunction gives a smaller variance on the energy estimator (Section 4.2.1). We further show in Section 4.2.2 that unlike a typical power method, we cannot expect that the walkers converge to the ground-state wavefunction as the variance of the individual walkers, controlled primarily by the chosen form of $B(\vec{x})$, is too large.
@@ 4.2.1. Convergence analysis of the energy
In this section, we analyze how the mixed energy estimator converges to the ground-state energy in one step of cp-AFQMC. We want to show that the variance of cp-AFQMC is controlled by $\left|c_{\perp}/c_{0}\right|$. When $\left|c_{\perp}/c_{0}\right|$ gets smaller, the variance of the energy gets smaller, i.e. the convergence of cp-AFQMC behaves like a vanilla power iteration. In an ideal case where $\left|c_{\perp}/c_{0}\right|=0$, i.e., the trial function is exactly the ground-state wavefunction $\Psi_{\mathrm{tr}}=\Psi_{0}$, there is a zero-variance principle. Thus, we are motivated to improve the trial wavefunction in Algorithm 2 to reduce the variance in ground-state energy estimation.
@@
Proof of both Lemma 4 and Theorem 5 can be found in the appendix. We notice that the energy convergence consists of a deterministic power iteration part and a fluctuating part that is roughly $a_{0}\|H-E_{0}I\|_{2}\tan\theta_{k}^{(n)}=\left|c_{\perp}/c_{0}\right|\|H-E_{0}I\|_{2}\tan\theta_{k}^{(n)}$ (when $a_{0}$ is small). Here $\|H-E_{0}I\|_{2}$ resembles the energy scale, and $\tan\theta_{k}^{(n)}$ is a random variable that captures the deviation of a random walker from the ground state. While currently, we do not have a characterization of the variance of $\tan\theta_{k}^{(n)}$, if we assume that it has a bounded variance of $O(1)$, when $a_{0}$ gets small, i.e. the trial function gets closer to the ground-state wavefunction, the fluctuation term can be made small. In an ideal case where $a_{0}=0$ so that the trial function is exactly the ground-state wavefunction $\Psi_{\mathrm{tr}}=\Psi_{0}$, there is no fluctuation in the energy estimation. As a matter of fact, we have
@@ 4.2.2. Convergence analysis of the walkers
where $\mathcal{P}_{\perp}$ is the orthogonal projector onto the space spanned by $\{\Psi_{1},\ldots\Psi_{2^{d}-1}\}$. We want to show in the following that $\tan\theta_{k}^{(n+1)}(\vec{x})<\tan\theta_{k}^{(n)}$ up to some noise terms. However, the noise terms will not vanish even for very good trial wavefunction.
@@
Proof of Theorem 6 can be found in the appendix. As we can see, even when $a_{0}=0$, i.e. the trial wavefunction is exactly the ground-state wavefunction, there is still an error term $O\left(\frac{\|\widetilde{B}_{k}^{(n)}(\vec{x})-e^{-\Delta\tau H}\|}{|\braket{\Psi_{0},\Phi_{k}^{(n)}}|}\right)$. Thus, the individual walker is not guaranteed to get closer to the ground state. In fact, in general, one needs an exponential number of walkers to get $\frac{\|\widetilde{B}_{k}^{(n)}(\vec{x})-e^{-\Delta\tau H}\|}{|\braket{\Psi_{0},\Phi_{k}^{(n)}}|}<1$. This is due to the fact that, generally, one needs to sample an exponential number of $\widetilde{B}_{k}^{(n)}(\vec{x})$ to approximate $\exp(-\Delta\tau H)$. In other words, when sampling a multi-dimensional function, we can evaluate various properties with a polynomial cost, but to map out the function to pre-specified details by relying on the sample positions can require exponential cost.